OpenStudy (anonymous):

Triangle ABC is an equilateral triangle with side lengths of 10 inches.... What is the length in inches, of AD?

OpenStudy (anonymous):

\[\sqrt{75}\]

OpenStudy (anonymous):

length of AD is given by 1/2 *sqrt(3) * length hence here AD =1/2 *sqrt(3) *10 = 5*sqrt(3)

OpenStudy (rajee_sam):

5 sqrt 3

OpenStudy (anonymous):

yup same as my answer